I need help with this algebra problem..
Solve by substitution
3x + 6y = 18
-6x + y = 16
Thank you for any help!!!

Answer by stanbon(75887)

Solve by substitution
3x + 6y = 18
-6x + y = 16
-----------------------------------------
Solve the 2nd equation for "y":
y = 6x+16
---
Substitute into the 1st equation:
3x + 6(6x+16) = 18
3x + 36x + 96 = 18
39x = -78
x = -2
----------------
Since y = 6x+16, y = 6*-2+16 = 4
----------------
Ans: (-2,4)
====================================
Cheers,
Stan H.